OpenStudy (anonymous):

(x+a)-(x+a)^3 step by !step please

OpenStudy (anonymous):

is it \[(x+a)-(x+a)^{3}\]

OpenStudy (anonymous):

factor out (x+a)

OpenStudy (anonymous):

you will get (x+a)(1-(x+a)^2)

OpenStudy (anonymous):

a^2-b^2 = (a+b)(a-b). Use that to solve the expression

OpenStudy (anonymous):

not to solve, i meant simplify the expression.

OpenStudy (anonymous):

did you understand, Lammy?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

what didnt you get?

OpenStudy (anonymous):

so far i have (x+a)^3 = (x+a)(x+a)^2 = (x+a)(x^2+2ax+a^2 correct?

myininaya (myininaya):

Does your problem say to factor?

OpenStudy (anonymous):

its cubing a binomial